OpenStudy (anonymous):

What is the 8th term of this geometric sequence? 2, 6, 18, 54…

OpenStudy (anonymous):

162

OpenStudy (anonymous):

is it correct?

OpenStudy (anonymous):

I wont know until finish the quiz, ill let you know

OpenStudy (anonymous):

Do u want the logic behind it?

OpenStudy (anonymous):

I would appreciate that

OpenStudy (anonymous):

1st no=2 2nd no=2*3=6 3rd no=6*3=18 4th no=18*3=54 5th no=54*3=112 got it??

OpenStudy (anonymous):

Ah ok, that makes sense, thank you!